OpenStudy (anonymous):

what is a number is divided by 5. Then 3 is added to the quotient. After subtracting 10, the result is 30. What is the number?

OpenStudy (anonymous):

30 = x/5 +3 -10

OpenStudy (anonymous):

so what is the number

OpenStudy (anonymous):

I don't give answers, I help in the best way I can. 30 = x/5 +3 -10 In order to isolate and solve for X, move the +3 to the left by subtracting 3 from both sides. Then move the -10 to the left by adding 10 to both sides. Then multiply both sides by 5 to get X.

OpenStudy (anonymous):

csso how do you set it up from the beginning so I can see all these steps please

OpenStudy (anonymous):

\[30 = \frac{ x }{ 5 } +3 -10\]

OpenStudy (anonymous):

isiso is 30 the numberor 10

OpenStudy (anonymous):

Problem: \[30=\frac{ x }{ 5 }+3−10\] Combine constants: \[30=\frac{ x }{ 5 }-7\] Isolate variable by moving constant to left. \[37=\frac{ x }{ 5 }\] Multiply both sides by 5 to get x.

OpenStudy (anonymous):

wwhat do you men by isolating constant to the left don't understand

OpenStudy (anonymous):

Kendra use 35 cents more in coupons at the store than Leanne. Leanne used 75 cents less thant tancy, who used 50 cent more thn Jaclyn. Jaclyn used 0 cents in coupons. what was the coupons Kendra used?

OpenStudy (anonymous):

Isolate the variable x: This means try to get x by itself. You have a constant -7 on the right. In order to get x more isolated, you have to move it to the other side by adding +7 to both sides.

OpenStudy (anonymous):

ok ok I understand Thanks now what about the other question I just posted